Please help.
Solve for k in k/k+1=3/k^2-k-2
Answer Options:
A. k=-1 and k=2
B. k=-1 and k=3
C. k=3
D. k=-1
Thank you!

Answer by rothauserc(4718)

Solve for k in k/k+1=3/k^2-k-2
factor denominator on right side of =
k /(k+1) = 3 / (k-2)*(k+1)
multiply both sides of = by (k-2)*(k+1)
k*(k-2) = 3
k^2 -2k -3 = 0
factor the above equation and we get
(k-3)*(k+1) = 0
to solve set (k-3) = 0 and (k+1) = 0
solutions are k = 3 and k = -1
and answer is B